OpenStudy (anonymous):

sketch of x^2 – 25y^2 = 25?

OpenStudy (anonymous):

divide by 25 both sides

OpenStudy (anonymous):

divide by 25

OpenStudy (anonymous):

so.... x^2- y^2=1

OpenStudy (anonymous):

\[\frac{x^{2}}{25}-\frac{y^{2}}{1}=1\] \[a^{2}=25 \] \[b^{2}=1 \]

OpenStudy (anonymous):

OH! Cool! thank you!